c is the midpoint of ad.b is the midpoint of ac. bc=12. what is the length of ad
1
Expert's answer
2020-11-01T17:27:22-0500


In the figure, C is the midpoint of AD.

"\therefore AC=CD" ...............(1)

B is the midpoint of AC.

"\therefore AB=BC" .................(2)

Given "BC=12" and we need to find "AD."

From (2), "AB=BC=12"

From figure, "AC=AB+BC=12+12=24"

From (1), "AC=CD=24"

From figure, "AD=AC+CD=24+24=48"

Thus, the length of AD is 48.